Let u=(2,6,-7), v=(-1,-1,8), and 3. If (2, 14, 11)=ku+lv what is the value of l?


1
Expert's answer
2021-08-09T15:34:17-0400

"\textsf{Given:}\\
u=(2,6,-7)\\
v=(-1,-1,8)\\
ku+lv= (2, 14, 11)----------(i)\\
ku=k(2,6,-7)=(2k,6k,-7k)\\
lv=l(-1,-1,8)=(-l,-l,8l)\\
ku+lv=(2k-l,6k-l,-7k+8l)---(ii)\\
\textsf{Comparing }(i)\textsf{ and }(ii),\\
2k-l=2\\
6k-l=14\\
-7k+8l=11\\
\textsf{Hence,}\\
2k-l-2=0------------(iii)\\
6k-l-14=0-----------(iv)\\
-7k+8l-11=0----------(v)\\
\textsf{By solving either (iii) and (iv), (iv) and (v),} \\
\textsf{or (iii) and (v) simultaneously, we obtain}\\
k=-3, l=-4.\\
\textsf{Hence, the value of } l=-4."